OpenStudy (anonymous):

In 1939 or 1940, Emanuel Zacchini took his human-cannonball act to an extreme. After being shot from a cannon, he soared over three Ferris wheels and into a net. Assume that he is launched with a speed of 26.5 m/s and at an angle of 53.0° and neglect air drag. a) Treating him as a particle, calculate his clearance over the first wheel. b) If he reach maximum height over the middle wheel, by how much did he clear it? c) How far from the cannon should the net's center have been positioned?

OpenStudy (anonymous):

See attached for accompanying picture.

OpenStudy (anonymous):

sooo what's the question? since the guy is being launched at an angle, break that up into x and y components and break up the initial velocity into x and y components as well

OpenStudy (anonymous):

x coordinate motion has no acceleration thus its \[x=v_{ox}t\] y coordinate motion has an acceleration thus you need to use \[y=-\frac{1}{2}at^2+v_{oy}t+x_{oy}\]

OpenStudy (anonymous):

@completeidiot I retyped the question to include ... the question. Lol... I've tried doing it but my answers come up different from the real answers, apparently: a) 5.3m b) 20m c) 69m

OpenStudy (anonymous):

gimme a sec

OpenStudy (anonymous):

Isn't b associated with the equation Hmax = \[\frac{ v _{0^{2}}\sin2\theta }{ 2g } \]

OpenStudy (anonymous):

I tried plugging in but my answer is always off

OpenStudy (anonymous):

dunno, i always just solve for the time it reaches max height using derivatives, then plugging that back into the height equation, then subtracting the height of wheel

OpenStudy (anonymous):

lemme work it out

OpenStudy (anonymous):

I know derivatives but we're just working with trig here so if I theoretically didn't know calc...

OpenStudy (anonymous):

meh fine -b/2a = vertex

OpenStudy (anonymous):

well thats the t coordiate so plug that back into heigh equation to find actual height, then subtract height of wheel

OpenStudy (anonymous):

 v_y=26.5 sin(53)=21.163 v_x=26.5 cos(53)=15.948 screw significant figures x_o =0 because you're landing on a net also 3 m tall y=-.5(9.8)t^2+ 21.163*t y=0=-4.9t+21.163 t=4.31 vx*4.31= total distance travelled=68.88m clearing the first whell 15.948m=vx determine time when he reaches 23 meters horizontally 23/15.948=1.44218 sec plug that in for t and determine height h(1.44218)=20.329 determine vertex cheating by using a graphing calculator t=2.1594s h=22.85m using the time value of the vertex, determine horizontal distance travelled 34.438m away from cannon

OpenStudy (anonymous):

oh wait you dont need to know how far away the middle wheel was

OpenStudy (anonymous):

... slightly confused but thank you

OpenStudy (anonymous):

also did you take into account that the initial vertical displacement was 3m from ground level @completeidiot

OpenStudy (anonymous):

you have a vertical displacement of 3, but the net is also a vertical displacement of 3 3-3 = 0

OpenStudy (anonymous):

you're landing on the net, not hte ground

OpenStudy (anonymous):

oh, right

OpenStudy (anonymous):

it would be quite painful to hit the ground at about 26.5m/s

OpenStudy (anonymous):

actually it would be faster, but only by a little

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

let me try and absorb this... (I got put into AP Physics C when I wanted Physics B, lol) a medal for your efforts

OpenStudy (anonymous):

if you dont understand anything, feel free to ask actually send me a message instead

OpenStudy (anonymous):

for b - I was given an angle so I guess I'm expected to use it... I would think it would be [(26.5)^2*sin2(53)]/2g but then I get a height of 34.4 which, minus 18, is 16.4m and not 20m

OpenStudy (anonymous):

whatever, I'll just put that down. I'm pretty tired